Propane reacts with oxygen to make carbon dioxide and water. If 150g of propane and 150g of oxygen produces 160g of CO2 and water, how many kilograms of propane and oxygen must be used to produce 1 kilogram of water?
1
Expert's answer
2020-11-11T10:35:18-0500

C3H8+ 5O2"\rightarrow" 3CO2 + 4H2O


according to stoichiometry...

1 mole of propane reacts with 5 moles of oxygen to produce 3 moles of CO2 and 4 moles of water.


moles of water = 1000/18 = 55.55moles


moles of propane required = 55.55/4 = 13.89 moles

mass of propane required = 13.89"\times" 44 = 611.66 grams


moles of oxygen required = 55.55"\times4" /5 = 44.44 moles

mass of oxygen required = 44.44"\times" 32 = 1422.08 grams